# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v2

@@ II.3 Equivalent definitions
We define $\mathbf{x_{1}}\equiv\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}$ and $\mathbf{y_{1}}\equiv\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}}$, and observe that ($\mathbf{x_{1}},\mathbf{y_{1}})$ lies within the infinitesimal neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ given that $\epsilon_{x}$ and $\epsilon_{y}$ are sufficiently small. Therefore, relabeling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at Eq. (14) exactly, thereby demonstrating that Definition 3a implies Definition 3c. Since these steps done to reach at this implication can be reversed, Definition 3c implies Definition 3a. Thus, we have established the equivalence between both these definitions. Moreover, in the light of the already established equivalence of Definitions 3a and Definition 3b, the equivalence of Definition 3b and Definition 3c is also automatically established.
It, however, is interesting to note that and-, sum-, and or-definitions are not equivalent and need not imply each other except that the and-definition implies the sum-definition for any ESS. This aspect of the definitions will be taken up in the next two sections.
@@ III Bimatrix games
Among the three sets of definitions which one is the most useful one? Or each of definitions works equally well? While there might be various arguments in favour of each one, one desirable aspect of ESS is probably undeniable: ESS is a static game-theoretic concept—it basically is about existence of an equilibrium solution; how one reaches it starting from a scenario that is not in equilibrium requires invoking some sort of dynamics. In evolutionary game theory, replicator equation Taylor and Jonker (1978); Cressman and Tao (2014) is the most accepted paradigm in this respect. Thus, one expects the replicator equation’s fixed points that correspond to ESS to be asymptotically stable, that is, we expect the following claim to be satisfied an ESS Cressman and Tao (2014); Cressman (1996):
Theorem 1: An ESS is a locally asymptotically stable fixed point of the following set of equations called replicator equations,
@@ III.1 Strong Stability
We note that while $x_{i}$ is the frequency of $i$th type in population X and $y_{i}$ is the frequency of $i$th type in population Y, we never explicitly specified what the corresponding strategies are for the two types. So let us bring it into consideration. So, let $m$-(pheno) types of players in population X with frequencies $\{x_{1},x_{2},\cdots,x_{m}\}(\equiv\mathbf{x}$) be, respectively, mapped to strategies $\{\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}\}$; similarly, we have $n$-(pheno) types in population Y with frequencies $\{y_{1},y_{2},\cdots,y_{n}\}$($\equiv\mathbf{y}$) be mapped to strategies $\{\mathbf{q}_{1},\mathbf{q}_{2},\cdots,\mathbf{q}_{n}\}$, respectively. In general, $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ are, respectively, elements of an $N_{1}$ dimensional simplex and $N_{2}$ dimensional simplex; there are two corresponding payoff matrices denoted as ${\sf U}_{1}$ and ${\sf U}_{2}$, respectively, representing the payoffs for role X and role Y players, should they adopt pure strategies in the respective simplices. The size of ${\sf U}_{1}$ is $N_{1}\times N_{2}$, while ${\sf U}_{2}$ has dimensions $N_{2}\times N_{1}$. In these notations, it is obvious that the matrices, ${\sf B}$ and ${\sf C}$, defined in the previous section are the ones with elements $b_{ij}=\mathbf{p}_{i}\cdot{\sf U}_{1}\mathbf{q}_{j}$ and $c_{ij}=\mathbf{q}_{i}\cdot{\sf U}_{2}\mathbf{p}_{j}$, respectively.
What we want to establish can be most clearly done in the simplest non-trivial setting where $m=n=2$. In any case, we would say (cf. Hofbauer and Sigmund (1998)), a strategy profile, $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, is strongly stable if a mean population strategy profile, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, evolves under the replicator equation to converge at $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. What is desired is that a strongly stable strategy be equivalent to evolutionarily stable strategy. Since the replicator equation is written using states and not strategies, we must first establish that the ESS conditions obeyed by strategies and states are equivalent.
@@
${\sf B}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{11}&c_{12}\\ c_{21}&c_{22}\end{pmatrix}.$ (18)
Suppose, without any loss of generality, $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is an ESS as per Definition 1b. It is trivial to see that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\implies b_{11}\geq b_{21}\implies\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}\geq\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}\implies c_{11}>c_{21}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}\geq\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. Furthermore in case the equalities hold in the relation between terms containing states, then it is easily seen that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\implies b_{12}>b_{22}$, which in turn means $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}$ when $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}=\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$; similarly, we can show $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}$ when $\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}=\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. In conclusion, if Definition 1b is satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$. It is rather repetitive to show that the conclusion remain unchanged had $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$—been chosen an ESS as per Definition 1b. Furthermore, employing these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b the state-strategy equivalence in the definition of respective ESS’s remains intact.
@@
We now effectively have to show that if an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then it is approached by neighbourhood initial conditions—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$. This, in fact, is the case can be argued as follows: Here $(x,y)=(0,0)$ denotes the ESS. We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and should those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are deciding factor for the stability of $(0,0)$. Consequently, comparing the dominating terms, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, which is the case of strict NE. However, (b) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ does not ensure that all the trajectories starting in the neighbourhood of $(0,0)$ would flow towards $(0,0)$—some flows to other phase points on the line $x=0$. Hence, local asymptotic stability of $(0,0)$ is not guaranteed; the fixed point is at best Lyapunov stable, a weaker form of stability. Similar ‘weak’ stability can be seen for the remaining two cases: (c) if $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$; and (d) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$. This weak stability criteria is equivalent to—as one can easily see from state-strategy correspondence discussed above—condition (ii) and (iii) in Definition 1b (and-definition) which is, thus, aptly also known as weak ESS Hofbauer and Sigmund (1998).
In summary, ESS under any of the three definitions is strongly stable if it is a strict NE, since the mean population strategy converges towards $(\mathbf{p}_{2},\mathbf{q}_{2})$. However, aforementioned weak stability can only be guaranteed by the and-definition.
@@ III.2.1 Strict NE as ESS
One can actually find that and-, sum-, and or- definitions are equivalent in the case ESS is strict NE; and linear stability analysis shows that satisfying any of these definitions guarantees the local asymptotic stability of the corresponding fixed point. However, since this exercise will be extended to more general asymmetric games—where both inter- and intra-specific interactions are considered—in Section IV.1, we refrain from providing the proof right here to avoid repetition. The argument presented in Section IV.1 can be easily adapted to the case of bimatrix games.
@@
Realizing that the two sums in $V_{1}$ represent Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, evidently $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$. If we additionally can show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})$ in some neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotically stable fixed point. To this end, on taking the time derivative, we get,
@@
where we have used Eq. (17). Of course, $\dot{V_{1}}<0$ when either inequalities in (5) or inequality in (10) is satisfied. Thus, any pure ESS that is an isolated fixed point, satisfying either Definition 1c or Definition 2c has $V_{1}$ as a Lyapunov function, ensuring that the corresponding fixed point is locally asymptotically stable. We mention that while a Lyapunov function for 2ESS (or-definition) is not known, the fixed point corresponding to the pure 2ESS can be shown to be locally asymptotically stable through alternate means (see Cressman (1996)).

### v3

@@ II.3 Equivalent definitions
Relabelling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at (14) exactly, thereby demonstrating that Definition 3a implies Definition 3c. Since these steps done to reach at this implication can be reversed, Definition 3c implies Definition 3a. Thus, we have established the equivalence between both these definitions. Moreover, in the light of the already established equivalence of Definitions 3a and Definition 3b, the equivalence of Definition 3b and Definition 3c is also automatically established.
@@ III Bimatrix games
It is an interesting fact that and-, sum-, and or-definitions are not equivalent and need not imply each other in general. So, among the three sets of definitions which one is the most useful one? Or each of the definitions works equally well? While there might be various arguments in favour of each one, one desirable aspect of ESS is probably undeniable: ESS is a static game-theoretic concept—it basically is about existence of an equilibrium solution; how one reaches it starting from a scenario that is not in equilibrium requires invoking some sort of dynamics. In evolutionary game theory, replicator equation Taylor and Jonker (1978); Cressman and Tao (2014) is the most accepted paradigm in this respect. Thus, one expects the fixed points of the replicator equation that correspond to ESS to be asymptotically stable, that is, we expect the following  proposition (cf. Thereom 4 in Ref. Cressman and Tao (2014), Theorem 4.5 and 4.6 in Ref. Cressman (1996)) to be satisfied an ESS:
Proposition 1: An ESS is a locally asymptotically stable fixed point of the following set of equations called the replicator equations,
@@ III.1 Strong Stability
We note that while $x_{i}$ is the frequency of $i$th type in population X and $y_{i}$ is the frequency of $i$th type in population Y, we never explicitly specified what the corresponding strategies are for the two types. So let us bring it into consideration. Let there be $m$-(pheno) types of players in population X with frequencies $\{x_{1},x_{2},\dots,x_{m}\}(\equiv\mathbf{x})$, which are mapped to strategies $\{\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{m}\}$, respectively. Similarly, let there be $n$-(pheno) types in population Y with frequencies $\{y_{1},y_{2},\dots,y_{n}\}(\equiv\mathbf{y})$, mapped to strategies $\{\mathbf{q}_{1},\mathbf{q}_{2},\dots,\mathbf{q}_{n}\}$, respectively. In general, $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ are, respectively, elements of an $N_{1}$ dimensional simplex and $N_{2}$ dimensional simplex; there are two corresponding payoff matrices denoted by ${\sf U}_{1}$ and ${\sf U}_{2}$, respectively, representing the payoffs for role X and role Y players, should they adopt pure strategies in the respective simplices. The size of ${\sf U}_{1}$ is $N_{1}\times N_{2}$, while ${\sf U}_{2}$ has dimensions $N_{2}\times N_{1}$. In these notations, it is obvious that the matrices, ${\sf B}$ and ${\sf C}$, defined in the previous section are the ones with elements $b_{ij}=\mathbf{p}_{i}\cdot{\sf U}_{1}\mathbf{q}_{j}$ and $c_{ij}=\mathbf{q}_{i}\cdot{\sf U}_{2}\mathbf{p}_{j}$, respectively.
What we want to establish can be most clearly done in the simplest non-trivial setting where $m=n=2$. In any case, we say (cf. Hofbauer and Sigmund (1998)), a strategy profile, $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, is strongly stable if, whenever $\mathbf{\hat{p}}$ is a convex combination of $\mathbf{p}_{1},\cdots,\mathbf{p}_{m}$ and $\mathbf{\hat{q}}$ is a convex combination of $\mathbf{q}_{1},\cdots,\mathbf{q}_{n}$, the mean population strategy profile, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, converges under the replicator equation to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. What is desired is a strongly stable strategy to be equivalent to evolutionarily stable strategy. Since the replicator equation is written using states and not strategies, we must first establish that the ESS conditions obeyed by strategies and states are equivalent.
@@
${\sf B}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}\leavevmode\nobreak\ \textrm{and}\leavevmode\nobreak\ {\sf C}=\begin{pmatrix}c_{11}&c_{12}\\ c_{21}&c_{22}\end{pmatrix}.$ (18)
Suppose, without any loss of generality, $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is an ESS as per Definition 1b. It is trivial to see that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\implies b_{11}\geq b_{21}\implies\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}\geq\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}\implies c_{11}{\color[rgb]{0,0,0}\geq}c_{21}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}\geq\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. Furthermore in case the equalities hold in the relation between terms containing states, then it is easily seen that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\implies b_{12}>b_{22}$, which in turn means $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}$ when $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}=\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$; similarly, we can show $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}$ when $\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}=\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. In conclusion, if Definition 1b is satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$. It is rather repetitive to show that the conclusion remain unchanged if $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$is chosen to be an ESS as per Definition 1b. Furthermore, employing these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b, the state-strategy equivalence in the definition of respective ESSes remains intact.
@@
We now have to show that if an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then it is approached by neighbourhood initial conditions—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$. This can be argued as follows: Here $(x,y)=(0,0)$ denotes the ESS. We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and if those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are the deciding factor for the stability of $(0,0)$. Consequently, comparing the dominating terms, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, which is the case of strict NE. However, (b) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ does not ensure that all the trajectories starting in the neighbourhood of $(0,0)$ would flow towards $(0,0)$—some flows to other phase points on the line $x=0$. Hence, local asymptotic stability of $(0,0)$ is not guaranteed; the fixed point is at best Lyapunov stable, a weaker form of stability. Similar ‘weak’ stability can be seen for the remaining two cases: (c) if $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$; and (d) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$. This weak stability criteria is equivalent to—as one can easily see from state-strategy correspondence discussed above—condition (ii) and (iii) in Definition 1b (and-definition) which is, thus, also known as weak ESS Hofbauer and Sigmund (1998).
In summary, ESS under any of the three definitions is strongly stable only if it is a strict NE, since the mean population strategy converges towards $(\mathbf{p}_{2},\mathbf{q}_{2})$. However, aforementioned weak stability can only be guaranteed by the and-definition.
@@ III.2.1 Strict NE as ESS
One can actually find that and-, sum-, and or- definitions are equivalent in the case ESS is strict NE; and linear stability analysis shows that being a strict NE guarantees the local asymptotic stability of the corresponding fixed point. However, since this exercise will be extended to more general asymmetric games—where both inter- and intra-specific interactions are considered—in Section IV.1, we refrain from providing the proof right here to avoid repetition. The argument presented in Section IV.1 can be easily adapted to the case of bimatrix games.
@@
Realizing that the two sums in $V_{1}$ represent the Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and the KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, evidently $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$. If we can additionally show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})$ in some neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotically stable fixed point. To this end, on taking the time derivative, we get,
@@
where we have used Eq. (17). Of course, $\dot{V_{1}}<0$ when either inequalities in (5) or inequality in (10) is satisfied. Thus, any pure ESS that is an isolated fixed point, satisfying either Definition 1c or Definition 2c has $V_{1}$ as a Lyapunov function, ensuring that the corresponding fixed point is locally asymptotically stable. We mention that while a Lyapunov function for 2ESS (or-definition) is not known, the fixed point corresponding to the pure 2ESS can be shown to be locally asymptotically stable through alternative means (see Ref. Cressman (1996)).